OpenStudy (anonymous):

If Log"2n"(1944) = Log"n" (486√2), determine the value of n^6. does anyone no how to do this?

OpenStudy (anonymous):

the "2n" and "n" are subscripts of log

OpenStudy (anonymous):

let Log"2n"(1944) = x then (2n)^x = 1944 also Log"n" (486√2) = x and n ^x = 486√2 (2n)^x / n ^x = 1944 / 486√2 2^x = 2.828427125 x ln 2 = ln 2.828427125 x = 1.5

OpenStudy (anonymous):

plug x=1.5 into n ^x = 486√2 n ^ 1.5 = 486√2 1.5 ln n = ln 486√2 n = 77.88148 n^6 = 2.2315 * 10^11 not sure if this is right ?- such a big number.

OpenStudy (anonymous):

i cant see any mistake after reviewing it again romoor though there might be another easier way of doing it

OpenStudy (anonymous):

lets see if the value of x and n satisfy the first log: (2 * 77.88148)^1.5 = 1944 - yes that's ok